OpenStudy (anonymous):

Factor completely 18a^2+24ad+8d^2

OpenStudy (anonymous):

I know 18x8 =144 and a sum of 24 what times what will give me 144 and equal 24

OpenStudy (anonymous):

2(9x^2+12x+4) =2(3x+2)^2

OpenStudy (anonymous):

I need two numbers

OpenStudy (anonymous):

oh..wait

OpenStudy (anonymous):

i need two numbers that will times to get 144 and =24

OpenStudy (anonymous):

and i will try and finsih solving equation

OpenStudy (anonymous):

2 is common factor, can be taken out

OpenStudy (anonymous):

12*12=144 12+12=24.if u wanna go with ur method

OpenStudy (anonymous):

0therwise =2(3a+2d)^2

OpenStudy (anonymous):

so that is the answer?

OpenStudy (anonymous):

yes